Please don’t forget to do part D and part E!! There are two firms.
Suppose that the market demand curve for mineral water is given as Q-100-10P and marginal cost is fixed at $4. Find the equilibrium price and quantity in each type of different market structure. Show your calculation (2 points for each subquestion). a) Monopoly b) Coumot duopoly c) Stackelberg duopoly d) Bertrand duopoly (MR is fixed at the level of MC). e) Perfect competitive market (MR is fixed at the level of MC)
Answer #1

Under monopoly equilibrium is attained where MR = MC.

Under cournot model, firms compete on quantities so quantity is split into Q1 and Q2.

Under Stackelberg model, one firm is follower and other is leader.

Under Bertrand model, firms compete on prices so price splits into P1 and P2.

Under perfect competition, equilibrium is where P = MC.

a) Monuo IOP = 100-6 LU 2. or2 usunat mco 13 O. 2 o, 2 Cesctrand Model qMRmc [00 - 00-to P, - 2 02- - P. =96-10A 96-10x3.2 До p = 3.2 c) Stackebe model Pasit (b) eqh 2. 2.ut Value Ove eauation 6 6-o.aq,ー3+ o. 1 a. =ο 30 uim 100-6 100-9 = yo